The following two linear functions represent a market (thus one is a supply function, the other a demand function). Circle the answer closest to being correct. Approximately what will suppliers willingly supply if the government controls the market price to be $3.00 (You must first find the market equilibrium price and quantity in order to see how the $3.00 relates to them)?

Q = 100 – 4.6P and Q = 75 + 6.2P

A. 2.3 B. 84.3 C. 86.2 D. 89.3 E. 93.1 F. 93.6 (all close, but approximate)
Answer #1

When price=3 then quantity demanded=100-4.6(3)=86.2

and quantity supplied=75+6.2(3)=93.6

Thus there is an excess supply of 7.4 and 86.2 quantity would be traded but supply is 93.6

ans is F